OpenStudy (anonymous):

Which expression is the simplified form of the square root of k to the twenty-first power.? A:k^5 B:k^10the square root of k. c:k^10 D:k^4the square root of k.

OpenStudy (helder_edwin):

\[ \sqrt{k}^21=\sqrt{k}\cdot\sqrt{k}^20=\sqrt{k}\cdot k^10 \] do you get it?

OpenStudy (helder_edwin):

sorry \[ \sqrt{k}\cdot k^{10} \]

OpenStudy (anonymous):

ok yeahh now i do. I was just a lil bit unsure but i get it. Thank you(: